OpenStudy (anonymous):

AB has endpoints A(3,6) and B(9, -2) and midpoint M. What is AM?

hero (hero):

First find M: \(M = \left(\dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2}\right)\)

OpenStudy (anonymous):

M= (6,2) right?

hero (hero):

In this case \((x_1,y_1) = (3,6)\) \((x_2, y_2) = (9,-2)\)

hero (hero):

To find the length of \(\overline{AM}\), use the distance formula: \(AM = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\)

hero (hero):

In this case \((x_1, y_1) = (3,6)\) \((x_2, y_2) = (6,2)\)

OpenStudy (anonymous):

does it equal 5?

hero (hero):

Correct

OpenStudy (anonymous):

thank you

hero (hero):

yw